### v4

# Survival Modeling from Whole Slide Images via Patch-Level Graph Clustering and Mixture Density Experts

Abstract
We present a modular framework for predicting cancer-specific survival from whole-slide pathology images (WSIs). The approach integrates four main components: (i) Quantile-Based Patch Filtering, which employs quantile-based thresholding to identify prognostically informative tissue regions; (ii) Graph Regularized Patch Clustering using $k$-NN graph to model phenotype-level heterogeneity through spatial–morphological coherence; (iii) Hierarchical Feature Aggregation for learning intra- and inter-cluster dependencies; and (iv) an Expert-Guided Mixture Density Modeling module to estimate complex survival distributions using Gaussian distributions. The proposed model achieves a concordance index of $0.653{\pm}0.037$ on TCGA-LUAD, $0.719{\pm}0.011$ on TCGA-KIRC, and $0.733{\pm}0.037$ on TCGA-BRCA, surpassing current state-of-the-art methods.
Index Terms—  Survival Analysis, Whole-Slide Images, Attention, Mixture Density Expert, Deep Learning.
1 Introduction
Precise survival prediction from histopathology images is essential for personalized oncology, enabling effective treatment planning and risk assessment. Whole-slide images (WSIs) encapsulate detailed tissue morphology—covering tumor structure and its microenvironmental context—but their enormous resolution and absence of localized labels pose challenges for direct modeling. Consequently, patch-based multiple instance learning (MIL) has emerged as the prevailing approach, segmenting each slide into smaller tiles for weakly supervised analysis.
Although MIL has proven effective for cancer grading and subtyping, survival prediction further requires modeling long-range dependencies and subtle morphological cues across distant tissue regions. Traditional methods like the Cox proportional hazards model fail to capture such complex, non-linear relationships, while deep learning models often lack interpretability and spatial awareness.
To overcome these challenges, we introduce a unified framework comprising four key modules: (1) Quantile-Based Patch Filtering, which adaptively selects prognostically informative tissue regions via quantile thresholding; (2) Graph-Regularized Patch Clustering, designed to capture spatial–morphological coherence among relevant patches; (3) Hierarchical Feature Aggregation, enabling the learning of intra- and inter-cluster dependencies; and (4) an Expert-Guided Mixture Density Modeling component for estimating complex survival distributions. Collectively, these modules facilitate interpretable and accurate survival prediction by integrating spatial organization, phenotypic representation, and probabilistic outcome modeling.
2 Related Work
Weakly supervised learning plays a pivotal role in WSI-based survival prediction, primarily utilizing multiple instance learning (MIL). Early approaches such as ABMIL [ilse2018attentionbaseddeepmultipleinstance], DSMIL [li2021dualstreammultipleinstancelearning], and CLAM [lu2020dataefficientweaklysupervised] employed attention pooling to identify prognostic patches but neglected spatial dependencies. Graph-based models like PatchGCN [chen2021slideimages2dpoint] incorporated spatial relationships, while hierarchical transformer models such as HIPT [chen2022scalingvisiontransformersgigapixel], HGT [10.1007/978-3-031-43987-2_72], and TransMIL [shao2021transmiltransformerbasedcorrelated] captured multi-scale representations at significant computational cost. More recent multimodal frameworks—HiLa-X [CuiJia_HiLa_MICCAI2025], PathoGen-X [10981028], and CoC [ZhoHai_CoC_MICCAI2025]—combine histopathology with textual, genomics, or methylation data modalities, while SCMIL [Yang_2024] models survival outcomes using a Normal distribution. Nevertheless, these methods lack explicit mechanisms to disentangle latent phenotypes or leverage structural coherence. Our expert-guided mixture density module bridges this gap by coupling phenotype-aware clustering with expert-based survival modeling, where specialized experts capture distinct prognostic cues and a gating network adaptively integrates their outputs for enhanced interpretability.
Fig. 1: (a) Proposed survival model. (b) Expert-guided Mixture Density Modeling block.
3 Methodology
[…]
$\mathbf{P}_{\text{feat}}\in\mathrm{R}^{n\times d}$ (1)
Here, $n$ represents the total number of extracted patches, and $d$ denotes the dimensionality of each feature vector.
The overall framework consists of four core modules: Quantile-Based Patch Filtering (QPF), Graph-Regularized Patch Clustering (GRPC), Hierarchical Feature Aggregation (HFA), and Expert-Guided Mixture Density Modeling (EGMDM).
3.1 Quantile-Based Patch Filtering (QPF)
To emphasize diagnostically significant regions, a learnable quantile-based Multi-Layer Perceptron (MLP) identifies task-relevant patches based on importance scores computed from model logits. For each WSI, a threshold $\tau_{q}$ corresponding to the $q$-quantile of logits is computed, preserving the top $(1-q)\times 100\%$ of patches.
$\mathcal{P}_{\text{sel}}=\{\mathbf{P}_{i}\mid\text{logit}_{i}>\tau_{q}\},\quad\mathcal{P}_{\text{rem}}=\{\mathbf{P}_{i}\mid\text{logit}_{i}\leq\tau_{q}\}.$ (2)
This dynamic filtering removes irrelevant patches, enabling subsequent modules to concentrate on semantically rich regions relevant to survival outcomes.
3.2 Graph-Regularized Patch Clustering (GRPC)
The selected patches $\mathcal{P}_{\text{sel}}\in\mathrm{R}^{m\times d}$ are clustered for both morphological and spatial consistency. Features and coordinates are normalized to a common scale, and cosine similarity measures morphological relationships:
$S_{\text{morph}}(i,j)=\frac{\langle\mathbf{P}_{i},\mathbf{P}_{j}\rangle}{\|\mathbf{P}_{i}\|\,\|\mathbf{P}_{j}\|},$ (3)
Meanwhile, spatial similarity is measured using an exponential kernel applied to the Euclidean distances $D_{ij}$ computed between the normalized patch coordinates:
$S_{\text{spatial}}(i,j)=\exp\!\left(-\frac{D_{ij}}{\sigma_{D}}\right),\quad\text{where }\sigma_{D}=\text{std}(D)+\varepsilon.$ (4)
A unified similarity matrix combines both cues:
$S=\omega_{\text{morph}}S_{\text{morph}}+\omega_{\text{spatial}}S_{\text{spatial}}$ (5)
Using the similarity matrix $S$, a $k$-nearest-neighbor (k-NN) graph is constructed, and GPU-accelerated K-Means guided by the graph topology clusters patches into $K$ groups $L_{1},\dots,L_{K}$}. Each cluster reflects a specific tissue microenvironment—like tumor epithelium, stroma, or necrosis—providing a biologically meaningful and spatially consistent structure within the slide.
3.3 Hierarchical Feature Aggregation (HFA)
To model dependencies within clusters, a two-level attention mechanism is employed.
Intra-cluster attention. Each cluster $L_{i}$ is processed using multi-head self-attention (MHSA) [vaswani2017attention] to capture fine-grained local relationships among patches:
$L^{\prime}_{i}=\text{LayerNorm}\!\big(L_{i}+\text{MHSA}(L_{i})\big).$ (6)
This step improves contextual understanding within morphologically similar regions.
Inter-cluster attention. The refined clusters are condensed into representative embeddings.
$R_{i}=\frac{1}{|L^{\prime}_{i}|}\sum_{\mathbf{x}\in L^{\prime}_{i}}\mathbf{x},$ (7)
which are subsequently passed through another MHSA layer to capture global relationships across regions:
$R^{\prime}=\text{LayerNorm}\!\big(R+\text{MHSA}(R)\big).$ (8)
Intra-cluster features are concatenated as $\widetilde{P}=\mathrm{Concat}(L^{\prime}_{k})_{k=1}^{K}$. Averaging $R^{\prime}$ across clusters produces a broadcast descriptor $R^{\prime}_{\mathrm{exp}}$, expanded to match $\widetilde{P}$ and added residually to yield $\widehat{P}$. Finally, $\widehat{P}$ is concatenated with unselected irrelavant patch features to form $\mathcal{P}_{\text{final}}$.
WSI-level aggregation. Slide-level representation is obtained through attention pooling as in AMIL [ilse2018attentionbaseddeepmultipleinstance].
$\mathbf{z}_{\text{WSI}}$ $=\sum_{i}\alpha_{i}\,\mathcal{P}_{\text{final},i},$ (9)
$\alpha_{i}$ $=\text{softmax}\!\big(W_{a}\,\tanh(W_{h}\,\mathcal{P}_{\text{final},i}^{\top})\big),$
where $\mathbf{z}_{\text{WSI}}$ represents the aggregated prognostic feature of the WSI.
3.4 Expert-Guided Mixture Density Modeling
An Expert-Guided Mixture Density Modeling block (Figure 1b) is utilized to estimate continuous survival probabilities by modeling the complete survival time distribution. Based on the mixture-of-experts paradigm [bishop2006pattern] and SurvMDN [han2022survivalmixturedensitynetworks], it employs multiple experts to construct Gaussian mixtures that represent diverse survival patterns.
The survival time $y$ corresponding to a whole slide image is modeled on its features $\mathbf{z}_{\text{WSI}}$, as illustrated in Figure 1b. Each expert $e$ within the mixture outputs a Gaussian Mixture Model (GMM) over $y$:
$p(y\mid\mathbf{z}_{\text{WSI}},e)=\sum_{k=1}^{K}\lambda^{(e)}_{k}(\mathbf{z}_{\text{WSI}})\,\mathcal{N}\!\big(y\mid\mu^{(e)}_{k},\sigma^{(e)2}_{k}\big),$ (10)
where each $k$ component follows a Gaussian distribution with mean $\mu^{(e)}_{k}$ and standard deviation $\sigma^{(e)}_{k}$, while mixture weights $\lambda^{(e)}_{k}$ are computed via softmax over the $K$ components. Learnable anchors ($(P_{\mu},P_{\sigma})\in\mathrm{R}^{K}$) represent shared mean and standard deviation vectors, linearly transformed for each expert:
$\mu^{(e)}=W^{(e)}_{\mu}P_{\mu},\qquad\sigma^{(e)}=\text{softplus}\!\big(W^{(e)}_{\sigma}P_{\sigma}\big),$ (11)
The survival time $y$ is mapped to an unconstrained variable $t$ via a softplus transformation $t=\log(1+e^{y})$, which ensures non-negativity in its Gaussian domain and produces the Jacobian factor $\frac{dy}{dt}=\frac{e^{t}}{e^{t}-1}$. Finally, a gating network $G(\mathbf{z}_{\text{WSI}})$ assigns softmax probabilities across experts, yielding the overall survival probability functions:
$\text{TPDF}(t\mid\mathbf{z}_{\text{WSI}})$ $=\frac{dy}{dt}\sum_{e,k}G_{e}(\mathbf{z}_{\text{WSI}})\,\lambda^{(e)}_{k}\,$
$\quad\times\mathcal{N}\!\big(g^{-1}(t)\mid\mu^{(e)}_{k},\sigma^{(e)2}_{k}\big)$ (12)
$\text{SPF}(t\mid\mathbf{z}_{\text{WSI}})=1-\sum_{e,k}G_{e}(\mathbf{z}_{\text{WSI}})\,\lambda^{(e)}_{k}\,\Phi\!\left(\frac{g^{-1}(t)-\mu^{(e)}_{k}}{\sigma^{(e)}_{k}}\right)$
where $\Phi$ is derived by integrating $\mathcal{N}$, TPDF denotes the transformed probability density of death, and SPF represents the survival probability at time $t$.
3.5 Training Objective and Regularization
The model is trained using a negative log-likelihood loss ($\mathcal{L}_{\text{NLL}}$) that accounts for both censored ($c=0$ i.e. event not yet occurred) and uncensored ($c=1$, i.e. event occurred) cases. Given an observed time $t_{d}$ and censoring indicator $c\in{0,1}$,
$\mathcal{L}_{\text{NLL}}=-\,c\,\log\!\big(\text{TPDF}(t_{d}\mid\mathbf{z}_{\text{WSI}})\big)-(1-c)\,\log\!\big(\text{SPF}(t_{d}\mid\mathbf{z}_{\text{WSI}})\big)$
To promote interpretability and specialization, a gating entropy loss is introduced ($\mathcal{L}_{\text{ent}}$).
$\mathcal{L}_{\text{ent}}=-{\textstyle\sum_{e}}\,G_{e}(\mathbf{z}_{\text{WSI}})\,\log\!\big(G_{e}(\mathbf{z}_{\text{WSI}})\big)$ (15)
The total loss is:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{NLL}}+\lambda_{\text{ent}}\,\mathcal{L}_{\text{ent}}.$ (16)
Fig. 2: Kaplan–Meier survival plots illustrating distinct high and low-risk group separation across TCGA cohorts.
Table 1: Left: TCGA results (TDC, higher is better). Right: Ablations on $q$ in QPF, GRPC, and HFA. Best TDC values in bold.
(a) TCGA results. Modality:I=image,G=genomic,M=methylation,L=language prompt.
Method Modality LUAD KIRC BRCA MEAN
AMIL [ilse2018attentionbaseddeepmultipleinstance] I $0.624\pm 0.038$ $0.638\pm 0.065$ $0.654\pm 0.013$ $0.639$
CLAM [lu2020dataefficientweaklysupervised] I $0.605\pm 0.051$ $0.676\pm 0.032$ $0.632\pm 0.045$ $0.638$
DSMIL [li2021dualstreammultipleinstancelearning] I $0.593\pm 0.065$ $0.655\pm 0.031$ $0.618\pm 0.021$ $0.622$
TransMIL [shao2021transmiltransformerbasedcorrelated] I $0.545\pm 0.037$ $0.632\pm 0.036$ $0.591\pm 0.013$ $0.589$
SetMIL [zhao2022setmil] I $0.620\pm 0.027$ $0.694\pm 0.052$ $0.635\pm 0.027$ $0.650$
PatchGCN [chen2021slideimages2dpoint] I $0.587\pm 0.019$ $0.675\pm 0.045$ $0.582\pm 0.037$ $0.615$
HIPT [chen2022scalingvisiontransformersgigapixel] I $0.549\pm 0.025$ $0.640\pm 0.037$ $0.625\pm 0.046$ $0.605$
HGT [10.1007/978-3-031-43987-2_72] I $0.607\pm 0.058$ $0.648\pm 0.018$ $0.638\pm 0.022$ $0.631$
SCMIL [Yang_2024] I $0.622\pm 0.015$ $0.688\pm 0.037$ $0.674\pm 0.048$ $0.661$
PathoGen-X [10981028] I+G $0.620\pm 0.008$ --- $0.670\pm 0.020$ $0.645$
CoC [ZhoHai_CoC_MICCAI2025] I+M --- $0.709\pm 0.048$ $0.654\pm 0.036$ $0.682$
HiLa [CuiJia_HiLa_MICCAI2025] I+L $0.643\pm 0.055$ --- $0.659\pm 0.044$ $0.651$
Ours I 0.653 $\pm$ 0.037 0.719 $\pm$ 0.011 0.733 $\pm$ 0.037 0.700
(b) Ablations.
Variant LUAD KIRC BRCA
[…]
4 Experimental Setup
4.1 Datasets
We assess the proposed model on three TCGA cohorts: Lung Adenocarcinoma (LUAD; 459 WSIs), Kidney Renal Clear Cell Carcinoma (KIRC; 509 WSIs), and Breast Invasive Carcinoma (BRCA; 956 WSIs) [tcga2013pan]. Each cohort comprises WSIs annotated with overall survival time, recorded in months from diagnosis to death (uncensored, c = 1) or until the last follow-up (censored, c = 0).
4.2 Training Configuration
The quantile of pacth filtering module is fixed at $q=0.25$, selecting the top $75\%$ of patches by their logits. In the k-NN graph, each patch links to 10 neighbors, which are grouped into $C$ clusters and processed through an 8-head hierarchical attention module. The Expert-Guided Mixture Density block employs five experts, each fitting a Gaussian Mixture Model (GMM) with $K=100$ components. Training runs for 20 epochs using Adam (lr = $2\times 10^{-4}$, weight decay = $1\times 10^{-3}$, dropout = 0.1) with a batch size of 1. Results are reported as mean $\pm$ standard deviation of the Time-Dependent Concordance Index (TDC) [han2022survivalmixturedensitynetworks] metric across 5-fold cross-validation, using an NVIDIA A6000 GPU.
5 Results and Analysis
5.1 Quantitative Results
Table 1(a) presents the model’s performance across all datasets. Our framework consistently attains the highest mean TDC, reflecting superior discrimination and calibration. Histopathology-based multiple instance learning (MIL) models—AMIL [ilse2018attentionbaseddeepmultipleinstance], CLAM [lu2020dataefficientweaklysupervised], DSMIL [li2021dualstreammultipleinstancelearning], TransMIL [shao2021transmiltransformerbasedcorrelated], and SetMIL [zhao2022setmil]—served as the backbones for the expert-guided mixture density module, with results reported for comparison. Survival-oriented MIL approaches such as HIPT [chen2022scalingvisiontransformersgigapixel], Patch-GCN [10981028], HGT [10.1007/978-3-031-43987-2_72], and SCMIL [Yang_2024] act as benchmarks. As multimodal frameworks [10981028, ZhoHai_CoC_MICCAI2025, CuiJia_HiLa_MICCAI2025] lack open-source implementations, their reported results are drawn from the original publications. By combining quantile-based patch filtering, graph clustering, hierarchical aggregation, and expert-guided mixture density modeling—each expert capturing distinct morphological or contextual cues—our approach effectively identifies survival-relevant regions and achieves state-of-the-art results across all datasets, even surpassing multimodal counterparts while using only histopathology data.
5.2 Ablation Studies and Interpretability
Ablation experiments were conducted to assess the impact of Quantile-Gated Patch Filtering, Graph-Regularized Patch Clustering, and Hierarchical Feature Aggregation (Table 1(b)). Varying the quantile showed that q=0.25q{=}0.25 offers the best trade-off, preserving prognostic regions while removing noise. Eliminating any module reduces TDC, confirming their complementary roles in denoising and spatial reasoning. For interpretability, patients were stratified into high- and low-risk groups using predicted survival scores, and Kaplan–Meier curves were plotted (Fig. 2). The curves show clear survival separation (log-rank p=0.049p{=}0.049 for LUAD, p=0.047p{=}0.047 for KIRC, p=0.089p{=}0.089 for BRCA), demonstrating that the expert-guided design captures prognostic morphology. Dynamic patch selection and graph-aware attention further enhance high-risk region localization and interpretability.
6 Conclusion
We introduced a unified framework for WSI-based survival prediction that combines quantile-regularized patch filtering, graph-based patch clustering, hierarchical feature aggregation, and expert-guided mixture density modeling. By capturing local–global tissue dependencies through clustered attention and modeling survival distributions using mixture experts, the approach enhances both calibration and discrimination. Experiments on TCGA-LUAD, TCGA-KIRC, and TCGA-BRCA demonstrate consistent improvements in time-dependent concordance compared to pathology-based and multimodal baselines. Future directions include extending the framework to multimodal and uncertainty-aware patient-level survival prediction for improved clinical robustness.
7 Compliance with ethical standards
This study employed publicly available human data from [tcga2013pan]; thus, no separate ethical approval was required under the open-access license.
8 Acknowledgments
This study’s findings are based on data from the TCGA Research Network [tcga2013pan].

### v6

@@
 Although MIL has proven effective for cancer grading and subtyping, survival prediction further requires modeling long-range dependencies and subtle morphological cues across distant tissue regions.
-Traditional methods like the Cox proportional hazards model fail to capture such complex, non-linear relationships, while deep learning models often lack interpretability and spatial awareness.
+Traditional methods like the Cox proportional hazards model fail to capture such complex, non-linear relationships, while previous deep learning models often lack interpretability and spatial awareness.
 To overcome these challenges, we introduce a unified framework comprising four key modules: (1) Quantile-Based Patch Filtering, which adaptively selects prognostically informative tissue regions via quantile thresholding; (2) Graph-Regularized Patch Clustering, designed to capture spatial–morphological coherence among relevant patches; (3) Hierarchical Feature Aggregation, enabling the learning of intra- and inter-cluster dependencies; and (4) an Expert-Guided Mixture Density Modeling component for estimating complex survival distributions.
@@
 $S=\omega_{\text{morph}}S_{\text{morph}}+\omega_{\text{spatial}}S_{\text{spatial}}$ (5)
-Using the similarity matrix $S$, a $k$-nearest-neighbor (k-NN) graph is constructed, and GPU-accelerated K-Means guided by the graph topology clusters patches into $K$ groups $L_{1},\dots,L_{K}$}.
+Using the similarity matrix $S$, a $k$-nearest-neighbor (k-NN) graph is constructed, and GPU-accelerated K-Means guided by the graph topology clusters patches into $G$ groups $L_{1},\dots,L_{G}$}.
 Each cluster reflects a specific tissue microenvironment—like tumor epithelium, stroma, or necrosis—providing a biologically meaningful and spatially consistent structure within the slide.
@@
 $R^{\prime}=\text{LayerNorm}\!\big(R+\text{MHSA}(R)\big).$ (8)
-Intra-cluster features are concatenated as $\widetilde{P}=\mathrm{Concat}(L^{\prime}_{k})_{k=1}^{K}$.
+Intra-cluster features are concatenated as $\widetilde{P}=\mathrm{Concat}(L^{\prime}_{g})_{g=1}^{G}$.
 Averaging $R^{\prime}$ across clusters produces a broadcast descriptor $R^{\prime}_{\mathrm{exp}}$, expanded to match $\widetilde{P}$ and added residually to yield $\widehat{P}$.
@@
 3.4 Expert-Guided Mixture Density Modeling
-An Expert-Guided Mixture Density Modeling block (Figure 1b) is utilized to estimate continuous survival probabilities by modeling the complete survival time distribution.
-Based on the mixture-of-experts paradigm [bishop2006pattern] and SurvMDN [han2022survivalmixturedensitynetworks], it employs multiple experts to construct Gaussian mixtures that represent diverse survival patterns.
+The Expert-Guided Mixture Density Modeling block (Figure 1b) estimates continuous survival probabilities by modeling survival times, with each expert focusing on distinct WSI features $\mathbf{z}_{\text{WSI}}$.
+Following the mixture-of-experts paradigm [bishop2006pattern] and SurvMDN [han2022survivalmixturedensitynetworks], it constructs Gaussian mixtures to represent diverse survival patterns.
 The survival time $y$ corresponding to a whole slide image is modeled on its features $\mathbf{z}_{\text{WSI}}$, as illustrated in Figure 1b.
@@
 $\mathcal{L}_{\text{NLL}}=-\,c\,\log\!\big(\text{TPDF}(t_{d}\mid\mathbf{z}_{\text{WSI}})\big)-(1-c)\,\log\!\big(\text{SPF}(t_{d}\mid\mathbf{z}_{\text{WSI}})\big)$
-To promote interpretability and specialization, a gating entropy loss is introduced ($\mathcal{L}_{\text{ent}}$).
+To promote interpretability and specialization, a gating entropy loss ($\mathcal{L}_{\text{ent}}$) is introduced for diverse expert usage.
 $\mathcal{L}_{\text{ent}}=-{\textstyle\sum_{e}}\,G_{e}(\mathbf{z}_{\text{WSI}})\,\log\!\big(G_{e}(\mathbf{z}_{\text{WSI}})\big)$ (15)
@@
 Fig. 2: Kaplan–Meier survival plots illustrating distinct high and low-risk group separation across TCGA cohorts.
-Table 1: Left: TCGA results (TDC, higher is better). Right: Ablations on $q$ in QPF, GRPC, and HFA. Best TDC values in bold.
+Table 1: Left: TCGA results (TDC, higher is better). Right: Ablations on $q$ in QPF, GRPC, HFA and $e$ in EGMDM. Best TDC values in bold.
 (a) TCGA results. Modality:I=image,G=genomic,M=methylation,L=language prompt.
@@
 […]
+[…]
 4 Experimental Setup
@@
 4.2 Training Configuration
-The quantile of pacth filtering module is fixed at $q=0.25$, selecting the top $75\%$ of patches by their logits.
-In the k-NN graph, each patch links to 10 neighbors, which are grouped into $C$ clusters and processed through an 8-head hierarchical attention module.
-The Expert-Guided Mixture Density block employs five experts, each fitting a Gaussian Mixture Model (GMM) with $K=100$ components.
-Training runs for 20 epochs using Adam (lr = $2\times 10^{-4}$, weight decay = $1\times 10^{-3}$, dropout = 0.1) with a batch size of 1.
-Results are reported as mean $\pm$ standard deviation of the Time-Dependent Concordance Index (TDC) [han2022survivalmixturedensitynetworks] metric across 5-fold cross-validation, using an NVIDIA A6000 GPU.
+The quantile of patch filtering module is fixed at qq=0.25, selecting the top $75\%$ of patches by their logits.
+The k-NN graph built on similarity matrix $S$, connects each node to 10 neighbors, clustering patch indices into $G$ groups of 64 each and those corresponding patches are processed by an 8-head MHSA in the HFA module.
+The Expert-Guided Mixture Density block employs five experts, each modeling a K=100K\!=\!100 Gaussian mixture.
+Training runs for 20 epochs with Adam (lr $2\times 10^{-4}$, weight decay $1\times 10^{-3}$), batch size 1.
+Performance is reported as mean $\pm$ std of the Time-Dependent Concordance Index (TDC) [han2022survivalmixturedensitynetworks] over 5-fold cross-validation on an NVIDIA A6000 GPU.
 5 Results and Analysis
@@
 5.2 Ablation Studies and Interpretability
-Ablation experiments were conducted to assess the impact of Quantile-Gated Patch Filtering, Graph-Regularized Patch Clustering, and Hierarchical Feature Aggregation (Table 1(b)).
-Varying the quantile showed that q=0.25q{=}0.25 offers the best trade-off, preserving prognostic regions while removing noise.
-Eliminating any module reduces TDC, confirming their complementary roles in denoising and spatial reasoning.
-For interpretability, patients were stratified into high- and low-risk groups using predicted survival scores, and Kaplan–Meier curves were plotted (Fig.
-2).
-The curves show clear survival separation (log-rank p=0.049p{=}0.049 for LUAD, p=0.047p{=}0.047 for KIRC, p=0.089p{=}0.089 for BRCA), demonstrating that the expert-guided design captures prognostic morphology.
-Dynamic patch selection and graph-aware attention further enhance high-risk region localization and interpretability.
+Ablation experiments evaluated the effects of Quantile-Gated Patch Filtering, Graph-Regularized Patch Clustering, Hierarchical Feature Aggregation, and Expert-Guided Mixture Density Modeling (Table 1(b)).
+A quantile of q=0.25q{=}0.25 offered the best trade-off, retaining prognostic regions while suppressing noise.
+Removing any module reduced TDC, confirming their complementary roles in denoising and spatial reasoning.
+Single-expert models underperformed, whereas five experts improved results, highlighting the value of multi-expert modeling.
+For interpretability, patients were stratified into high- and low-risk groups using predicted survival scores.
+Kaplan–Meier curves (Fig.
+2) showed clear separation (log-rank p=0.049p{=}0.049 LUAD, p=0.047p{=}0.047 KIRC, p=0.089p{=}0.089 BRCA), validating prognostic relevance and regional localization.
 6 Conclusion